### v1

# Modeling Platelet P2Y1/12 Pathway to Integrin Activation

Abstract
Through experimental studies, many details of the pathway of integrin IIb3 activation by ADP during the platelet aggregation process have been mapped out. ADP binds to two separate G protein coupled receptors on platelet surfaces, leading to alterations in the regulation of the small GTPase RAP1. We seek to (1) gain insights into the relative contributions of both pathways to RAP1-mediated integrin activation and to (2) predict cell behavior in response to a continuous range of external agonist concentrations. To this end, we develop a dynamical systems model detailing the action of each protein in the two pathways up to the regulation of RAP1. We perform a parameter estimation using flow cytometry data to determine a number of unknown rate constants. We then validate with already published data; in particular, the model confirmed the effect of impaired P2Y1 receptor desensitization or reduced RASA3 expression on RAP1 activation. We then predict the effect of protein expression levels on integrin activation and show that components of the P2Y12 pathway are critical to the regulation of integrin. This model aids in our understanding of interindividual variability in platelet response to ADP and therapeutic P2Y12 inhibition. It also provides a more detailed view of platelet activation in the ongoing mathematical study of platelet aggregation.
[…]
A detailed dynamical systems model of integrin IIb3 activation mediated by ADP is presented in this report. This model takes in years of experimental literature on the relevant pathways into simulations that are efficient to run and simple to manipulate parameters to match a platelet mutation. Unlike in experiments, our model can finely adjust any parameter and analyze how a platelet responds to any number of experimental settings. We present the utility of such a model through parameter study on the copy number of proteins key to integrin activation and show that integrin response changes more significantly in response to changes in the P2Y12 pathway over the P2Y1 pathway.
Introduction
Platelets are the primary cellular component in arterial blood clots, supporting the coagulation response and recruitment of other cells/proteins to a growing thrombus. They perform a variety of processes that aid in clot growth and stabilization, such as releasing agonists into the environment to recruit other platelets, changing shape through cytoskeletal remodeling to increase surface area for reactions, and activating integrins on their surfaces for binding crosslinking proteins. For the purposes of this work, we will use the words “integrin activation” to refer to the conversion of integrin from a low fibrinogen affinity state to a high-affinity state.
Several molecules serve as receptor ligands that lead to integrin activation (1). Most ligands, like thrombin or collagen, initiate pathways that leads to irreversible activation of the small protein RAP1, which complexes with the integrin on its cytosolic end and leads to a conformation change on the extracellular end. Signaling pathways for other receptor-ligand bindings have been extensively studied by both experimentalists and modelers. Dunster et al. modeled the activation signals associated with collagen binding the GPVI receptor (2) and Lenoci et al. and Sveshnikova et al. modeled the irreversible activation of small GAP protein RAP1 by the PAR1 receptor (3, 4).
ADP binds to two different receptors, the G-protein coupled receptors (GPCR) P2Y1 and P2Y12, and is unique in that it leads to a partial, transient activation of integrin that lasts only a few minutes. Experimental work has uncovered much about the pathways that lead to the activation of integrin. Activation of P2Y1 induces activation of RAP1 via the calcium-dependent RAP-GEF, CalDAG-GEFI (CDGI), while activation of P2Y12 induces RAP1 activation through the inhibition of the RAP-GAP, RASA3 (5). The regulation of RAP1 is vital to maintaining platelet quiescence in the absence of clot formation, and therefore understanding the individual contribution of CalDAG-GEFI and RASA3 in integrin activation is of interest. Other models have been developed to study platelet activation and aggregation mediated by ADP (6), but to the authors’ knowledge, no detailed signaling pathways for ADP dependent activation have been developed.
In this work, we convert the collected information on both pathways to a system of ordinary differential equations (ODEs) that can be solved numerically to simulate a single platelet’s response to a given amount of P2Y1 and/or P2Y12 agonist. We use experimental data to estimate unknown parameters and validate the model against recently published data. We then explore how a simulated platelet responds to agonists under a change to the expression level of specified proteins. Unlike in experimental settings, we are able to tune expression with much finer precision to gain a more quantitative measure of sensitivity. We begin by describing the chemical signaling pathways upon platelet GPCR binding to ADP.
Biological Background
Figure 1 shows schematically how the P2Y1 and P2Y12 receptors are activated by ADP and, through different pathways, cause changes to both the activator and inhibitor of RAP1.
The extracellular domains of the platelet P2Y1 and P2Y12 receptors bind to ADP in the blood plasma. Upon binding to ADP, these receptors act as GEF enzymes on the membrane-associated G proteins Gq and Gi, respectively. The G proteins then unbind from the receptor and bind to membrane-associated enzymes PLC and PI3K, respectively. Both PLC and PI3K use the lipid PIP2 as substrate: PLC converts PIP2 to the membrane-associated molecule DAG and the cytosolic molecule IP3, and PI3K converts PIP2 to membrane-associated PIP3. These inositol lipids are part of a large cycle of formation and degradation. IP3 binds to IP3 receptors (IP3R) embedded in the dense tubular system (DTS) membrane and triggers the release of calcium ions from the DTS into the cytosolic space. Platelets contain other intracellular calcium stores, e.g., the mitochondrial and acidic stores (7), which impact calcium levels on a significantly longer timescale and therefore are ignored for this study. The chemically-gated IP3R channel is a four-subunit protein; each subunit contains an activating binding site for IP3, an activating binding site for calcium, and a slower inactivating binding site for calcium (8). Thus, the initial release of calcium from the DTS leads to a stage of positive feedback followed by negative feedback.
At rest, cytosolic calcium levels are maintained by exchanges with the plasma, in which calcium is assumed to have a concentration around 1 . The model incorporates the action of the PMCA pump, a passive leak current on the plasma membrane, and the action of the SERCA pump across the DTS membrane.
Calcium’s effect on integrin activation is mediated through its binding to the GEF enzyme CalDAG-GEFI. CalDAG-GEFI contains two EF domains which each binds to calcium with a dissociation constant (9). Calcium-bound CalDAG-GEFI converts RAP1 from its inactive GDP-bound form to its active GTP-bound form. The active form of RAP1 complexes with a variety of proteins on the cytosolic tail of , namely TALIN and KINDLIN3, to trigger integrin activation. RASA3 is a membrane-associated protein constitutively active in converting active RAP1 to its inactive GDP-bound form. RASA3 can be inactivated by interacting with PIP3, although the exact mechanism has not been determined (10). CalDAG-GEFI and RASA3 have nearly equal copy numbers in mouse platelets (approximately 30,000 ) whereas RAP1 in its two major isoforms has a total copy number of approximately 200,000 (11).
Materials and Methods
We used Mass Action or Michalis-Menten kinetics to describe each reaction in the P2Y1 and P2Y12 pathways. The concentration, denoted by , or surface density, denoted by , of each type of molecule in its various states was tracked using an ordinary differential equation (ODE). The overall model, therefore, comprised an extensive coupled system of ODEs that we must solve simultaneously, and we accomplished this using LLNL’s DLSODE solver package (12).
Here, we discuss the setup of the differential equation system. We used the model of Purvis et al. (13) as a starting point for our modeling; below, we discuss explicitly where we deviated and extended their model. The system was described by 82 differential equations; see the supplemental material for a complete list of the model’s equations. We then describe the statistical methods used to inform unknown model parameters. Finally, we describe the experimental methods used to generate the data against which we validate the model.
Model components and numerical setup
ADP Receptors:
Following prior modeling work on G protein coupled receptors (14), we explicitly tracked the state changes of the P2Y1 and P2Y12 receptors as they bind to ADP and their specific G protein. Each binding, unbinding, and catalytic reaction is described using the Law of Mass Action. Receptors bound to an inactive G protein can activate it, regardless of whether ADP is also bound to the receptor. However, the presence of ADP significantly increases the receptor’s efficiency.
We made two notable changes to the model as implemented in Purvis et al. (13). First, this and other previous models include an "inactive" GPCR state converted into an "active" state through a reversible, unimolecular reaction. We noted through numerical simulations (not shown) that the concentration for each inactive state quickly reached an equilibrium value proportional to the concentration for the corresponding active state. Therefore, we performed a quasi-steady-state (QSS) reduction to reduce the number of explicit states.
Second, we included a PKC-dependent inactivated state for P2Y1. We assumed that ADP-bound P2Y1 can be inactivated via phosphorylation of an intracellular site and reactivated via dephosphorylation. PKC mediates the phosphorylation process. The kinetic parameters for this reaction were estimated in this study.
Second Messengers:
The model explicitly tracked the second messenger proteins PLC and PI3K in their inactive state and active (i.e., bound to G protein) state (15). Binding and unbinding of G proteins were tracked through Mass Action kinetics, with an additional irreversible hydrolysis term, in which the second messenger both converts the subunit to a GDP bound form and releases it into the cytosol. Once activated, PLC and PI3K act upon the phosphoinositol species PIP2, converting it to IP3 and PIP3, respectively. These inositol species are recycled through a variety of states by various other phosphatases and kinases (16, 17, 18, 19, 20, 21, 22, 23).
Calcium Dynamics:
The flux of calcium ions across passive elements (namely, the IP3R and PM leak) are governed by the Nernst Equation. For example, the rate at which calcium moves across the DTS membrane through conducting IP3R channels is given by:
where is the copy number of IP3R subunits, is the probability a single channel is in a conducting state, is the conductance of the channel, is the universal gas constant, is the temperature, is the charge of a calcium ion, is Faraday’s constant, and and are the free calcium concentrations in the DTS and cytosol, respectively. We model the IP3R channel using the six-state model given by Sneyd and Dufour (8) and use their formula for computing , which is dependent on the fraction of receptors in the two conducting states.
Once calcium is released into the cytosol, it can bind to various pumps, enzymes, and buffers. We modeled both the SERCA and PMCA pumps using 6-state transport models (24), where two calcium ions at a time are transported across their respective membranes. We explicitly modeled PKC binding of calcium and DAG using Mass Action kinetics. As described above, PKC molecules bound to both calcium and DAG can inactivate the P2Y1 receptor. To match the experimental setup, we included the binding of calcium to a fluorescent experimental probe with a concentration of 5 and a dissociation constant of 100 . This probe was only included during the parameter estimation that relied on calcium data. In all other experiments, we excluded the probe.
GEF/GAP Enzymes:
We modeled both the binding of calcium with CalDAG-GEFI (9) and the binding of PIP3 with RASA3 using Mass Action kinetics. We then modeled the activation and inactivation of RAP1 using Michalis-Menten reaction terms, such that the differential equation for the surface density of RAP1-GTP is given by:
where is the surface area of the cell membrane, is the volume of the cytosol, and the catalytic rates and Michalis-Menten constants are estimated in this work.
Simulation Protocol:
The copy number of each protein within mouse platelets has been documented using proteomics (11). From the data, we computed the concentration or surface density of every species in the model. The model is run in two stages: an equilibrium stage and an experimental stage. In the equilibrium stage, we started by assuming that every molecule is in a form that is not bound to any other molecule, and we assumed there is no external ADP. We then ran the system to steady-state, which we determined by computing the relative change in protein states at each iteration and stopping when changes decreased below a threshold. The concentration and surface density values at steady state were then used as the initial conditions in the experimental stage, where at , the external ADP concentration was instantaneously increased to a set value, and the simulation was run for 600 s.
Sensitivity Analysis:
We briefly describe the Method of Morris, a one-at-a-time global sensitivity analysis algorithm. The algorithm gives a way to estimate the derivative of a model output with respect to model parameters when no closed form of the output exists and when dealing with many model parameters. In this study, we are interested in describing the derivative of the maximum RAP1-GTP value with respect to protein copy numbers.
The method begins by defining a hypercube in parameter space; for each dimension, the center is given by the average value of a specified protein’s copy number, and the half-width is given by the standard deviation of that protein’s copy number. Beginning at the center, the algorithm defines several paths to the boundary of the hypercube such that only one parameter changes with each step along a path. The paths were computed using Python’s SALib package (25, 26). Each vertex on a path specifies a parameter set used for a single simulation. In each simulation, we recorded the peak number of RAP1-GTP and then used finite differences to approximate the derivative of the maximum RAP1-GTP concentration with respect to the single parameter changed in that step. This process generated a sampling of the derivative of peak RAP1-GTP as a function of each parameter within the hypercube; for each parameter, we reported the mean and standard deviation of the approximation to the derivative. A large value of suggests a high sensitivity with respect to the given parameter, and a large value of suggests a high correlation between the given parameter and other parameters in the model. We again relied on SALib to compute and for each parameter.
Total integrin response metric:
JON/A is a fluorescent probe that binds tightly to the integrin in its high-affinity state. JON/A is assumed to bind irreversibly to the integrin, and enough JON/A is added to ensure rapid binding between the two species. To compare model outputs to previously published JON/A binding assays, we reported the integral with respect to time of RAP1-GTP over the 600 s after agonist application. See the supplemental for further details.
Experimental procedures and data extraction
Flow cytometry experiments:
Washed platelets in Tyrode’s buffer were diluted to platelets/mL and loaded with 5 M Fluo-4 (Thermo Fisher Scientific) for 30 minutes at 37∘C in the dark. Afterward, the samples were diluted to platelets/mL in Tyrode’s buffer and activated with 10 M ADP in the presence of 1 mM Ca2+ and Alexa647-labeled fibrinogen (100 g/ml, Sigma) while being continuously sampled on a BD C6 Plus flow cytometer. Kinetic calcium mobilization and fibrinogen binding were analyzed in Flow Jo (Version 10) as mean fluorescence intensities over time.
Parameter estimation:
Previous models (13) of calcium signaling in cells and experimental work (11) with mice platelets have provided many of the rate constants and concentrations needed for this study. However, the rates governing the GEF and GAP action on RAP1 have yet to be entirely determined, and parameters estimated to fit prior calcium data were inconsistent with ours. Therefore, we performed a parameter estimation study to compare model outputs to our experimental data given in Figure 2. We first describe the method of converting fluorescence intensity to a measure of concentration and then the parameter estimation method.
For each dataset provided, we recorded the maximum intensity at select times to generate a single representative curve for each experiment. To convert calcium probe fluorescence intensity to concentration, we assumed that a relative increase in fluorescent intensity corresponds to an equivalent relative increase in the protein of interest. For calcium, we assumed a resting concentration of 40 and a peak concentration of 200 , in line with the literature (1). This data was used to estimate kinetic parameters for P2Y1, , SERCA, and PMCA. To reduce the number of parameters to estimate, we make the simplifying assumption that the kinetic parameters for P2Y12, Gi, and PI3K are the same as those for P2Y1, Gq, and PLC.
For fibrinogen data, we performed a similar rescaling, then assumed 1) that there are no fibrinogen bound to resting platelets and 2) a WT platelet activated by a saturating level of ADP allows an a priori known quantity of integrins to activate and 3) fibrinogen binds instantly to such an integrin. Finally, we used a 1:1 stochiometric relationship between RAP1-GTP and integrin to arrive at the amount of RAP1-GTP that is active intracellularly. We assume that either a maximum of 1% or 10% of integrins become activated, which equated to a maximum of 500 or 5000 RAP1-GTP, respectively; for either assumption, we ran a separate parameter estimation for rates governing CalDAG-GEFI, RASA3, and RAP1.
For any choice of parameters, we ran the model according to the simulation protocol described above. From the list of model outputs, we take the concentration of free cytosolic calcium and the number of activated RAP1 proteins, then compute the least squared error between the model and the datasets at specified times. The goal of the parameter estimation algorithm is to minimize this error, and for this, we relied on MATLAB’s fmincon function.
Results and Discussion
Model outputs estimated to match experimental data:
Figure 3a compares a calcium spike dataset to the model’s calcium spike. Because the data and simulated calcium timecourse in the wildtype and RASA3 heterozygote are nearly identical, only the wildtype timecourse is given. The simulated calcium curves approximated the data well but smoothed over some of the dynamics near the peak.
Figures 3b-c show a comparison between the fibrinogen binding data and the simulated RAP1-GTP spike using the lower bound approximation (see Methods) of the data in Figure 3b and the upper bound approximation of the data in Figure 3c. For each of these assumptions about the number of integrins activated by ADP, our estimation procedure sought to minimize the sum of the mean squared differences between data and simulation for both the WT and RASA3+/- cases. Our model approximated the data well when ADP was applied to the peak RAP1-GTP response. However, simulations showed a slower return to baseline than the experimental data. For estimated parameter values, see Supplemental Tables LABEL:tab:CaEstimatedParams and LABEL:tab:RAP1EstimatedParams.
Model sensitivity to protein copy number is unaffected by our data approximation:
Many of the simulations conducted in this work involve adjusting protein copy numbers. Therefore, understanding how our assumptions about the copy number data affected model outputs was useful. Thus, the global sensitivity analysis algorithm, the Method of Morris, was used to determine how changes in protein copy numbers change the maximum amount of RAP1-GTP seen in simulations. All protein copy numbers in the model were selected from the experimental literature (11), and each copy number was increased or decreased by one standard deviation from the mean.
Figure 4 shows the sensitivity of the maximum RAP1 response to the protein copy numbers. Comparing the two subfigures, we see little difference in the sensitivity of the model to copy numbers. We also note that in both figures, the same proteins have the most sensitivity, namely Gi, P2Y12, PMCA, CalDAG-GEFI, and RASA3, meaning that a change in copy number of these proteins should yield similar relative responses. This served as evidence that under the appropriate normalization, the results generated by one estimated parameter set are similar to those generated by the other. We therefore present results using only our upper bound estimate parameter set for the remainder of this work.
Moreover, Figure 4 shows that the peak level of RAP1-GTP was most sensitive to changes in the P2Y12 receptor and Gi protein. This implies that the variability in expression of P2Y12 or Gi had the most significant effect on RAP1 activation. In particular, equivalent variability in the expression of RASA3, the immediate effector of RAP1, did not produce the same variation in the RAP1-GTP signal.
model predicts saturating integrin response to ADP:
We then examined how our simulated platelet responds to a range of applied agonist concentrations. Figure 5 shows how the numbers or concentrations of various chemicals changed over time for a given application of ADP at s. Figures 5a and 5c show that immediately upon application of ADP into the system, there was a rapid rise in the number of P2Y1 and P2Y12 receptors, respectively, bound with ADP. When 200 M of ADP was applied, essentially all P2Y1 and P2Y12 receptors were bound to ADP at the peak. Due to PKC’s desensitization of the P2Y1 receptor, the number of active P2Y1 receptors decreased from its peak value as the simulation progressed. Interestingly, the number of active P2Y1 receptors decreased most rapidly for the highest concentration of ADP. This decrease was matched by a fast rise in the number of desensitized P2Y1 receptors, as shown in Figure 5b. Therefore, as the concentration of ADP applied increased, the maximum response increased, but the time interval on which a near-peak level was maintained decreased.
Following receptor activation, second messengers PLC and PI3K became active by binding the appropriate G protein, leading to a rise in the production of the inositol species IP3 and PIP3, respectively. Figure 5d shows that the IP3 concentration rose with a slight delay compared to the receptor concentrations. Later in the simulations, as the P2Y1 receptors become desensitized, the rate of formation of IP3 decreased, leading to a decline in its concentration.
IP3 then activated IP3R, which led to an influx of calcium into the cytosol from the DTS. As seen in Figure 5f, the level of free cytosolic calcium increased, plateaued, and then decreased throughout the simulation. The rise in cytosolic calcium levels led to the activation of proteins PKC and CalDAG-GEFI, the latter of which is seen in Figure 5g following a similar pattern of rise, plateau, and decrease as the corresponding calcium trace.
Concurrently with IP3, calcium, and CalDAG-GEFI spiking, the P2Y12 pathway led to the formation of the PIP3 inositol species as seen in Figure 5e, followed by the inactivation of RASA3, as seen in Figure 5h. Unlike in the P2Y1 pathway, the lack of a desensitization mechanism for P2Y12 meant that PIP3 was constantly created by PI3K during the simulation, leading to a permanent decrease in the number of active RASA3 molecules.
The activation of CalDAG-GEFI and inactivation of RASA3 caused an increase in the amount of RAP1 bound to GTP, shown in Figure 5i. When the CalDAG-GEFI levels decreased, RAP1-GTP levels also decreased and approached their baseline levels. As shown more explicitly in Figure 8, the model’s response to ADP saturated around 10 M, so the 200 M ADP simulated experiment represents the maximum ADP response.
RAP1 activation response increases with applied agonist:
We next attempted to validate our model’s output by comparison to data previously published (27). To do this, we simulated platelets’ exposure to ADP, which binds to both P2Y1 and P2Y12 receptors, and platelets’ exposure to the synthetic agonist MRS2365, which binds only to the P2Y1 receptor. To accommodate the synthetic agonist, the model was changed in two ways: 1) by setting the binding rate between agonist and P2Y12 to zero, i.e., , and 2) by lowering the dissociation constant for the agonist binding to P2Y1 from 0.6 to 0.01 M, i.e. M.
Figure 6 shows how WT platelets respond to the synthetic agonist. We found that the application of a saturating amount of MRS2365 yielded similar levels of P2Y1 and CalDAG-GEFI activation compared to those evoked by the application of a saturating amount of ADP (compare Figure 6a with Figure 5a and Figure 6b with Figure 5g). However, because the concentration of RASA3 remained high throughout the MRS2365 simulations, RAP1 activation was greatly limited.
We next considered platelets with a mutation to the intracellular tail of their P2Y1 receptor, which makes it unable to be phosphorylated and therefore the receptor is prevented from being desensitized to agonists and being taken up into the cell (27). These platelets are known as P2Y platelets or simply 340-0P. Figure 7 shows how a 340-0P platelet responds to ADP. While the peak value of various proteins was similar to the wildtype case seen in Figure 5, the simulated mutant platelet’s P2Y1 signal never declined, and consequently the downstream calcium, CDGI, and RAP1 signals also did not decline.
To gauge variations in platelet responses due to variations in agonist concentration, we performed computational experiments for a wide range of applied ADP and MRS2365 concentrations. In the corresponding physical experiments, the readout was the amount of the integrin probe JON/A bound irreversibly to activated integrin. As a proxy for the amount of JON/A bound at time we used the time-integral of the RAP1-GTP curve up to that time. Figure 8 shows the results of these experiments. As expected, for both agonists, the RAP1-GTP response increased as the concentration of the agonist was increased. This response was consistently greater for 340-0P platelets than for WT ones. This is attributed to the fact that for the P2Y platelets, the P2Y1 receptors remained active and continued to perform their GEF action on the Gq protein for the duration of the simulation. Consequently, the IP3, calcium, and RAP1-GTP responses are prolonged, and therefore, the area under the RAP1-GTP curve increased.
Two differences were evident in how both wildtype and 340-0P platelets responded to the two stimuli. First, the response to MRS2365 began at lower concentrations than that to ADP, consistent with the higher affinity of MRS2365 for the P2Y1 receptor. Second, the response to a saturating concentration of MRS2365 was lower than for a saturating concentration of ADP, consistent with unabated RASA3 GAP activity in the MRS2365-stimulated platelets.
We compared the results of Figures 2c and 4c in (27), recreated in Figure 8a and 8b, to the results of Figures 8c and 8d. First, comparing the experiments in which platelets are stimulated with ADP in Figure 8a and 8c, we saw an increase in the integrin response over the same range of applied ADP concentration for both wildtype and P2Y platelets. Additionally, we saw that for a given applied ADP concentration, the 340-0P platelet signal was consistently larger than the wildtype platelet. We note that in Figure 8b and 8d, we saw a similar relationship between mutant and wildtype platelets. Deviations between the model and experiment are evident when comparing simulated versus experimental mutant platelets stimulated with MRS2365, where the experiments show a more gradual increase in integrin activation as a function of applied agonist concentration compared to simulations. In general, comparing our model to results from (27) showed good agreement, even when a single experimental modification is considered. As more modifications are combined into simulated platelets, the results may deviate from experiments.
Changes to protein copy number in the P2Y12 pathway affect RAP1-GTP response more than in the P2Y1 pathway:
The model allowed for the continuous variation of parameters, specifically protein copy numbers. We leveraged this to assess the effects of protein expression levels on the integrin activation response to a specified stimulus. In each set of simulations, we set the copy number of one of the proteins RASA3, CDGI, P2Y1, or P2Y12 to a value between 5% and 150% of its literature value (while holding other copy numbers at their literature values). We stimulated the platelet with 10 M ADP.
Figure 9 shows the results of varying the RASA3 or CDGI copy numbers. We saw that a decrease in the number of RASA3 molecules or an increase in the number of CDGI molecules elicits an increase in the RAP1-GTP response. Figure 9a shows that when the copy number of RASA3 is decreased by 50%, integrin activation was stronger and more prolonged than the same simulations with 100% RASA3 expression. In contrast, Figure 9b shows that increasing the amount of CDGI by 50% did not significantly change the peak level but did increase the length of the RAP1-GTP response.
Figure 9c plots the RAP1-GTP response as a function of a specified protein’s copy number, where we saw that the sensitivity of the response to RASA3 copy number variations was much greater than that for CDGI variations. Changes in the CDGI copy number led to a small and approximately linear change in RAP1 activation. In contrast, responses to decreases in the RASA3 copy number are nonlinear, with a sharp increase in sensitivity as the RASA3 copy number decreases to and beyond 50% of its baseline value.
Figure 10 shows a similar dichotomy between responses to variations in copy numbers of the ADP receptors P2Y1 and P2Y12. With P2Y12 at 150%, integrin activation was much greater at all times, and high levels were maintained for an extended period compared to 100%. With P2Y12 at 50%, integrin activation is weak. With P2Y1 at 50%, integrin activation was approximately half that of activation at 100%. There was only a small difference in integrin activation for P2Y1 100% and 150%. Examining model outputs, in this case, showed that although the amount of activated IP3 increased, the peak concentration of calcium did not increase significantly beyond 200 nM (data not shown).
Conclusion
In this report, we developed a mathematical model of integrin activation on a platelet surface and fitted it to experimental data. Thanks to years of experimental studies, a chemical pathway has been defined sufficiently to allow such a model to exist. Previous platelet activation models rely on other major indicators, such as calcium signals or aggregometry readings (28, 29). While effective, inferring integrin activation information from only one of the two ADP-dependent pathways does not fully encapsulate the regulatory pathway’s complexity, making it difficult to extend to particular mutant platelets. Our model can be used in testing and making predictions without the need for a large number of experimental trials. It is quite economical to run; for reference, a single simulation takes approximately one minute to run ten simulated minutes on a standard personal computer. Thus, data generation is significantly faster than generation of an equivalent amount of data in the lab.
While our estimated parameters yielded a model that matched our experimental dataset from the point of ADP addition up to the peak, our model tended to lengthen the duration of the RAP1-GTP signal. Under the assumption that our estimated parameters are the ideal parameters for the model, a mechanism for desensitization of some signaling protein in the P2Y12 pathway, leading to a recovery of RASA3 levels, could shorten the RAP1-GTP signal duration. Without sufficient evidence for a biological mechanism, we save modeling this portion of the pathway for future work.
In experiments where we varied the applied ADP concentration, we found that a platelet’s response saturates beyond 10 M, consistent with experiments. Our highly detailed model allowed us to examine the individual responses of each protein in time. We showed the usefulness of our model beyond directly comparing to experimental data by varying parameters that would be difficult to vary continuously in the lab, like protein expression levels, and making predictions on integrin activation. In particular, we showed that the total integrin response varies more significantly with changes to proteins in the P2Y12 pathway compared to the P2Y1 pathway.
This model can be used to conduct initial tests on the behavior of platelets under any form of experimental condition before running experiments. In addition, the outputs of this model describe integrin activation over time in far more detail than is typical in mathematical models for hemostasis and thrombosis (30, 31), where the tracking of platelet aggregation and fibrin formation is of interest. This may lead to improvements in these larger-scale models.
Author Contributions
Patel and Fogelson constructed the mathematical model and wrote the article. Patel carried out all simulations. Bergmeier conducted all experiments.
Acknowledgments
This work was supported by the National Institutes of Health, National Heart, Lung, and Blood Institute (grant R35 HL144976 to WB and R01 HL151984 to AF), as well as the National Science Foundation’s Graduate Research Fellowship to KP.

### v2

@@
 Abstract
-Through experimental studies, many details of the pathway of integrin IIb3 activation by ADP during the platelet aggregation process have been mapped out.
+Through experimental studies, many details of the pathway of integrin activation by ADP during the platelet aggregation process have been mapped out.
 ADP binds to two separate G protein coupled receptors on platelet surfaces, leading to alterations in the regulation of the small GTPase RAP1.
-We seek to (1) gain insights into the relative contributions of both pathways to RAP1-mediated integrin activation and to (2) predict cell behavior in response to a continuous range of external agonist concentrations.
+We seek to (1) gain insights into the relative contributions of both pathways to RAP1-mediated integrin activation and to (2) predict wildtype and mutated cell behavior in response to a continuous range of external agonist concentrations.
 To this end, we develop a dynamical systems model detailing the action of each protein in the two pathways up to the regulation of RAP1.
@@
 […]
-A detailed dynamical systems model of integrin IIb3 activation mediated by ADP is presented in this report.
+A detailed dynamical systems model of integrin activation mediated by ADP is presented.
 This model takes in years of experimental literature on the relevant pathways into simulations that are efficient to run and simple to manipulate parameters to match a platelet mutation.
-Unlike in experiments, our model can finely adjust any parameter and analyze how a platelet responds to any number of experimental settings.
-We present the utility of such a model through parameter study on the copy number of proteins key to integrin activation and show that integrin response changes more significantly in response to changes in the P2Y12 pathway over the P2Y1 pathway.
+Our model can finely adjust any parameter and analyze how a platelet responds to any number of experimental settings.
+We present the model’s utility through parameter study on the expression levels of proteins key to integrin activation and show that integrin response changes more significantly in response to changes in the P2Y12 pathway over the P2Y1 pathway.
+This insight can be used to understand how certain antithrombotic therapies, like clopidogrel, can affect a platelet’s intracellular signaling pathways.
 Introduction
@@
 Several molecules serve as receptor ligands that lead to integrin activation (1).
-Most ligands, like thrombin or collagen, initiate pathways that leads to irreversible activation of the small protein RAP1, which complexes with the integrin on its cytosolic end and leads to a conformation change on the extracellular end.
-Signaling pathways for other receptor-ligand bindings have been extensively studied by both experimentalists and modelers.
+Most ligands, like thrombin or collagen, initiate pathways that lead to irreversible activation of the integrin.
+Signaling pathways for these receptor-ligand bindings have been studied by both experimentalists and modelers.
 Dunster et al.
-modeled the activation signals associated with collagen binding the GPVI receptor (2) and Lenoci et al.
-and Sveshnikova et al.
-modeled the irreversible activation of small GAP protein RAP1 by the PAR1 receptor (3, 4).
-ADP binds to two different receptors, the G-protein coupled receptors (GPCR) P2Y1 and P2Y12, and is unique in that it leads to a partial, transient activation of integrin that lasts only a few minutes.
-Experimental work has uncovered much about the pathways that lead to the activation of integrin.
-Activation of P2Y1 induces activation of RAP1 via the calcium-dependent RAP-GEF, CalDAG-GEFI (CDGI), while activation of P2Y12 induces RAP1 activation through the inhibition of the RAP-GAP, RASA3 (5).
-The regulation of RAP1 is vital to maintaining platelet quiescence in the absence of clot formation, and therefore understanding the individual contribution of CalDAG-GEFI and RASA3 in integrin activation is of interest.
-Other models have been developed to study platelet activation and aggregation mediated by ADP (6), but to the authors’ knowledge, no detailed signaling pathways for ADP dependent activation have been developed.
+modeled the activation signals associated with collagen binding the GPVI receptor (2).
+Lenoci et al.
+(3) and Sveshnikova et al.
+(4) modeled platelet activation through the action of thrombin on the PAR1 receptor.
+We focused our modeling efforts on platelet inside-out integrin activation induced by ADP, an autocrine/paracrine small molecule agonist released by activated platelets.
+Our main reasons for choosing ADP as the agonist include (i) its physiological relevance as a key recruiter of free-flowing platelets to a growing aggregate in vivo and is the target of certain antithrombotic therapies such as clopidogrel (5), (ii) the well-defined signaling pathways connecting the ADP receptors, P2Y1 and P2Y12, to integrin activation mediated by the adapter protein, Talin1, and (iii) a relatively weak and reversible integrin activation response that provides a perfect system to model both positive and negative modulation of these signaling pathways.
+Furthermore, we focus specifically on the effect of inside-out signaling in our first version of this model.
+We experimentally isolate the effect of inside-out signaling by using the fluorescent molecule JON/A(PE) (6), which selectively binds to the active conformation of the integrin and prevents outside-in signaling.
+ADP binds to two platelet GPCRs, the Gq-coupled P2Y1 and Gi-coupled P2Y12 receptors.
+Activation of P2Y1 induces activation of phospholipase C (PLC) and formation of Ca2+ and diacylglycerol (DAG), while activation of P2Y12 induces PI3 kinase-dependent formation phosphatidylinositol-3,4,5-trisphosphate (PIP3).
+Ca2+ activates the RAP-GEF, CalDAG-GEFI (CDGI), while PIP3 inhibits the RAP-GAP, RASA3 (7).
+Activation of both pathways, but not P2Y1 or P2Y12 signaling alone, leads to robust and prolonged activation of small GTPase RAP1, a key regulatory component of the integrin activation complex (IAC) (8, 9, 10).
+The IAC consists of RAP1 and adapter proteins such as Talin1 (11).
+Following RAP1 activation, these proteins assemble at the intracellular tail of the integrin to induce a conformational change required for ligand binding.
+Once RAP1-GTP is converted to its GDP bound form, however, these proteins quickly unbind, and the integrin returns to its inactive conformation.
+Thus, the dynamics of RAP1 activation can be used as a proxy for integrin activation.
+It is also important to note that ADP-dependent mobilization of intracellular calcium stores is also linked to the influx of calcium from the extracellular space.
+However, this so-called store-operated calcium entry (SOCE) does not impact ADP-induced integrin inside-out activation (12).
+Furthermore, P2Y12 signaling leads to decreased levels of cyclic AMP, a second messenger that leads to the inhibition of various platelet activation responses (13).
+We have not included these pathways in our model as our previous work demonstrated that signals directly affecting CalDAG-GEFI and RASA3 function are most relevant for RAP1-dependent integrin inside-out activation.
+Other models (14) have been developed to study integrin activation and aggregation mediated by ADP, but to the authors’ knowledge, no detailed models of signaling pathways for ADP-dependent activation have been developed.
 In this work, we convert the collected information on both pathways to a system of ordinary differential equations (ODEs) that can be solved numerically to simulate a single platelet’s response to a given amount of P2Y1 and/or P2Y12 agonist.
 We use experimental data to estimate unknown parameters and validate the model against recently published data.
-We then explore how a simulated platelet responds to agonists under a change to the expression level of specified proteins.
+We then explore how a simulated platelet responds to agonists under a change to the expression level of P2Y1, P2Y12, RASA3, and CalDAG-GEFI.
 Unlike in experimental settings, we are able to tune expression with much finer precision to gain a more quantitative measure of sensitivity.
-We begin by describing the chemical signaling pathways upon platelet GPCR binding to ADP.
+We begin by describing the chemical signaling pathways triggered upon platelet GPCR binding to ADP.
 Biological Background
@@
 IP3 binds to IP3 receptors (IP3R) embedded in the dense tubular system (DTS) membrane and triggers the release of calcium ions from the DTS into the cytosolic space.
-Platelets contain other intracellular calcium stores, e.g., the mitochondrial and acidic stores (7), which impact calcium levels on a significantly longer timescale and therefore are ignored for this study.
 The chemically-gated IP3R channel is a four-subunit protein; each subunit contains an activating binding site for IP3, an activating binding site for calcium, and a slower inactivating binding site for calcium (15).
 Thus, the initial release of calcium from the DTS leads to a stage of positive feedback followed by negative feedback.
-At rest, cytosolic calcium levels are maintained by exchanges with the plasma, in which calcium is assumed to have a concentration around 1 .
-The model incorporates the action of the PMCA pump, a passive leak current on the plasma membrane, and the action of the SERCA pump across the DTS membrane.
+In resting platelets, the cytosolic Ca2+ concentration is 20-100 nM, a level that is maintained by active sequestration of Ca2+ in the DTS and the extracellular space.
+In addition to the IP3R channel, the model incorporates the active PMCA pump and passive leak current on the plasma membrane, and the active SERCA pump across the DTS membrane.
 Calcium’s effect on integrin activation is mediated through its binding to the GEF enzyme CalDAG-GEFI.
-CalDAG-GEFI contains two EF domains which each binds to calcium with a dissociation constant (9).
+CalDAG-GEFI contains two EF domains, which each binds calcium with a dissociation constant (16).
 Calcium-bound CalDAG-GEFI converts RAP1 from its inactive GDP-bound form to its active GTP-bound form.
-The active form of RAP1 complexes with a variety of proteins on the cytosolic tail of , namely TALIN and KINDLIN3, to trigger integrin activation.
+The active form of RAP1 complexes with a variety of proteins on the cytosolic tail of , for example, TALIN, to trigger integrin activation.
 RASA3 is a membrane-associated protein constitutively active in converting active RAP1 to its inactive GDP-bound form.
 RASA3 can be inactivated by interacting with PIP3, although the exact mechanism has not been determined (17).
-CalDAG-GEFI and RASA3 have nearly equal copy numbers in mouse platelets (approximately 30,000 ) whereas RAP1 in its two major isoforms has a total copy number of approximately 200,000 (11).
+CalDAG-GEFI and RASA3 have nearly equal copy numbers in mouse platelets (approximately 30,000 ), whereas RAP1 in its two major isoforms has a total copy number of approximately 200,000 (18).
 Materials and Methods
@@
 Once activated, PLC and PI3K act upon the phosphoinositol species PIP2, converting it to IP3 and PIP3, respectively.
-These inositol species are recycled through a variety of states by various other phosphatases and kinases (16, 17, 18, 19, 20, 21, 22, 23).
+These inositol species are recycled through a variety of states by various other phosphatases and kinases (23, 24, 25, 26, 27, 28, 29, 30).
 Calcium Dynamics:
@@
 Beginning at the center, the algorithm defines several paths to the boundary of the hypercube such that only one parameter changes with each step along a path.
-The paths were computed using Python’s SALib package (25, 26).
+The paths were computed using Python’s SALib package (32, 33).
 Each vertex on a path specifies a parameter set used for a single simulation.
@@
 However, simulations showed a slower return to baseline than the experimental data.
-For estimated parameter values, see Supplemental Tables LABEL:tab:CaEstimatedParams and LABEL:tab:RAP1EstimatedParams.
+For estimated parameter values, see Supplemental Tables S1 and S2.
 Model sensitivity to protein copy number is unaffected by our data approximation:
@@
 In the corresponding physical experiments, the readout was the amount of the integrin probe JON/A bound irreversibly to activated integrin.
-As a proxy for the amount of JON/A bound at time we used the time-integral of the RAP1-GTP curve up to that time.
+As a proxy for the amount of JON/A bound at time , we used the time-integral of the RAP1-GTP curve up to that time.
 Figure 8 shows the results of these experiments.
@@
 Deviations between the model and experiment are evident when comparing simulated versus experimental mutant platelets stimulated with MRS2365, where the experiments show a more gradual increase in integrin activation as a function of applied agonist concentration compared to simulations.
-In general, comparing our model to results from (27) showed good agreement, even when a single experimental modification is considered.
+Comparing our model to results from (10) showed good agreement, even when a single experimental modification is considered.
 As more modifications are combined into simulated platelets, the results may deviate from experiments.
@@
 The model allowed for the continuous variation of parameters, specifically protein copy numbers.
-We leveraged this to assess the effects of protein expression levels on the integrin activation response to a specified stimulus.
+We leveraged this to assess the effects of protein expression levels on the integrin activation response over a range of ADP stimuli.
 In each set of simulations, we set the copy number of one of the proteins RASA3, CDGI, P2Y1, or P2Y12 to a value between 5% and 150% of its literature value (while holding other copy numbers at their literature values).
-We stimulated the platelet with 10 M ADP.
 Figure 9 shows the results of varying the RASA3 or CDGI copy numbers.
@@
 In contrast, Figure 9b shows that increasing the amount of CDGI by 50% did not significantly change the peak level but did increase the length of the RAP1-GTP response.
-Figure 9c plots the RAP1-GTP response as a function of a specified protein’s copy number, where we saw that the sensitivity of the response to RASA3 copy number variations was much greater than that for CDGI variations.
-Changes in the CDGI copy number led to a small and approximately linear change in RAP1 activation.
-In contrast, responses to decreases in the RASA3 copy number are nonlinear, with a sharp increase in sensitivity as the RASA3 copy number decreases to and beyond 50% of its baseline value.
-Figure 10 shows a similar dichotomy between responses to variations in copy numbers of the ADP receptors P2Y1 and P2Y12.
-With P2Y12 at 150%, integrin activation was much greater at all times, and high levels were maintained for an extended period compared to 100%.
-With P2Y12 at 50%, integrin activation is weak.
-With P2Y1 at 50%, integrin activation was approximately half that of activation at 100%.
-There was only a small difference in integrin activation for P2Y1 100% and 150%.
-Examining model outputs, in this case, showed that although the amount of activated IP3 increased, the peak concentration of calcium did not increase significantly beyond 200 nM (data not shown).
+Figure 9c plots the RAP1-GTP response of platelets with specified expression level changes as a function of ADP concentration.
+We found that as we vary the expression of RASA3, the dose-response curve deviates further from the WT curve than for equivalent variations in the expression of CDGI.
+Furthermore, the variation in dose-response occurs across applied ADP concentrations.
+For example, at 10 nM ADP, there are modest increases in the integrin response as we vary expression levels within 50% of baseline and much more dramatic increases as we decrease RASA3 expression down to 5%.
+This result mirrors behaviors seen experimentally in Stefanini et al.
+(7), where a specific mouse platelet mutation resulting in 5% RASA3 expression led to an increase in both the baseline and stimulated integrin response.
+When we examined the integrin response as a function of expression level, we found that the dose-response curve scales approximately linearly with changes in CDGI expression (Supplemental Figure S1b).
+In contrast, platelet response as a function of RASA3 expression is highly nonlinear, with a dramatic dose-response curve as the RASA3 copy number decreases to and below 50% of its baseline value.
+We next were interested in determining how platelets would respond to ADP when their receptor expression levels varied.
+In Figure 10, we vary the expression of P2Y1 and P2Y12 individually to see the platelet’s response to ADP over time and the total dose response.
+We expected that the integrin response should be positively correlated with the expression levels of either receptor, and Figures 10a-b agree.
+Examining model outputs further verifies that the calcium timecourse is affected by variation in P2Y1 expression levels but not by variation in P2Y12 expression levels (data not shown).
+Figure 10c shows the dose-response curves for different receptor expression levels.
+Similar to RASA3 in Figure S1b, we found that variation in P2Y12 yielded larger changes to the dose-response curve than in P2Y1.
+Interestingly, Figures 10b-c show that an increase in the P2Y1 receptor to 150% leads to an almost negligible change in integrin response.
+Further analysis of model outputs shows that although increasing P2Y1 expression led to an increase in peak IP3 concentrations, the peak concentration of calcium did not increase significantly beyond 200 nM (data not shown).
+We show in Supplemental Figure S2a that the relative integrin response between 100% and 150% P2Y1 expression is more significant when 0.5 M ADP is used due to a difference in peak cytosolic calcium levels.
+Similar results were seen experimentally in Heckler et al.
+(34), in which they overexpress P2Y1 by 84% and notice an increase in aggregation response to low doses of ADP.
 Conclusion
-In this report, we developed a mathematical model of integrin activation on a platelet surface and fitted it to experimental data.
-Thanks to years of experimental studies, a chemical pathway has been defined sufficiently to allow such a model to exist.
-Previous platelet activation models rely on other major indicators, such as calcium signals or aggregometry readings (28, 29).
+In this report, we developed a mathematical model of inside-out integrin activation on a platelet surface and fitted it to experimental data.
+Thanks to years of experimental studies (1), a chemical pathway has been defined sufficiently to allow such a model to exist.
+Previous platelet activation models rely on other major indicators, such as calcium signals or aggregometry readings (35, 36).
 While effective, inferring integrin activation information from only one of the two ADP-dependent pathways does not fully encapsulate the regulatory pathway’s complexity, making it difficult to extend to particular mutant platelets.
@@
 Thus, data generation is significantly faster than generation of an equivalent amount of data in the lab.
+The results in this work rely on a combination of kinetic rates found in the literature or estimated from data.
+Copy numbers of each protein were based on proteomics analysis of mouse platelets (18).
+In some cases, due to a lack of available data, rate constants associated with a protein isoform found in human platelets were used (23, 25, 26, 28, 30).
+This choice allows us to estimate fewer parameters under the assumption that the action of the protein isoform found in mouse platelets acts similarly on its substrate compared to these human platelet isoforms.
 While our estimated parameters yielded a model that matched our experimental dataset from the point of ADP addition up to the peak, our model tended to lengthen the duration of the RAP1-GTP signal.
@@
 In particular, we showed that the total integrin response varies more significantly with changes to proteins in the P2Y12 pathway compared to the P2Y1 pathway.
+By decreasing the expression level of RASA3 to well below 50%, we found that both a resting and stimulated platelet presented a marked increase in integrin response, similar to results seen in Stefanini et al.
+(7).
+By changing expression levels of the receptors, we show that the integrin response can be modulated only above 0.1 M ADP.
+In Supplemental Figure S2a, we show that when using 50 nM ADP, we start to see a change in the integrin response between 100% and 150% expression of P2Y1, akin to the results of Hechler et al (34) in which they overexpress P2Y1.
+The results from this study can be used to understand how certain pharmacological agents affect the integrin activation process.
+In particular, the P2Y12 antagonist clopidogrel is taken long-term as an antithrombotic treatment, and its action is analogous to a reduction in the overall expression of P2Y12.
+More detailed modeling is needed to show this connection explicitly, and will be the subject of future work.
 This model can be used to conduct initial tests on the behavior of platelets under any form of experimental condition before running experiments.
-In addition, the outputs of this model describe integrin activation over time in far more detail than is typical in mathematical models for hemostasis and thrombosis (30, 31), where the tracking of platelet aggregation and fibrin formation is of interest.
+In addition, the outputs of this model describe integrin activation over time in far more detail than is typical in mathematical models for hemostasis and thrombosis (37, 38), where the tracking of platelet aggregation and fibrin formation is of interest.
 This may lead to improvements in these larger-scale models.
@@
 Patel and Fogelson constructed the mathematical model and wrote the article. Patel carried out all simulations. Bergmeier conducted all experiments.
+Declaration of Interests
+The authors declare no competing interests.
 Acknowledgments